Amirah and Anusha make a total of 560 toys in the ratio 12:16.
a)Find the number of toys Anusha makes.
b)the amount of money Amirah earns if she is paid PKR 165 for each toy.

Found 2 solutions by Theo, greenestamps:
Answer by Theo(13342)

x = number of toys that amirah makes.
y = number of toys that anusha makes.
x + y = 560
x/y = 12/16
solve for x to get x = 12/16 * y = 3/4 * y
x + y = 560 becomes 3/4 * y + y = 560
combine like terms to get 7/4 * y = 560
solve for y to get y = 560 * 4/7 = 80 * 4 = 320
this makes x = 560 - 320 = 240
amirah makes 240 toys
anusha makes 320 toys
at PKR 165 per toy, amirah makes 165 * 240 = PKR 39600.

Answer by greenestamps(13198)



Simplify the given ratio; 12:16 = 3:4

That ratio means Anusha makes 4/(4+3) = 4/7 of the toys.

a) The number of toys that Anusha makes is (4/7)*560 = 4*80 = 320.

ANSWER: 320

b) Anusha makes 320 toys, so Amirah makes 560-320 = 240 toys. If she is paid PKR 165 for each toy, then the amount she earns is PKR (165*240)= PKR 39600.

ANSWER: PKR 39600